OpenStudy (anonymous):

A silverwire has a resistance of 1.5Ω at 0°C and a temperature coefficient of resistance of 0.00375/°C. To what temperature must be the raised to double the resistance?

OpenStudy (anonymous):

this is not physics section

OpenStudy (ash2326):

@alfers101 This is a Physics question. Please post Physics questions in Physics Subject. Thanks Nevertheless I'll help you here :)

OpenStudy (ash2326):

@alfers101 are you here?

OpenStudy (anonymous):

yup i am. im just desperate to answer this.

OpenStudy (ash2326):

Cool:D We are given that the resistance of the silver ware at 0 °C is 1.5 ohm Temperature coefficient of resistance is 0.00375 ohms/° C so for each 1° increase in temp. resistance will increase by 0.00375 ohms We want to find the temp at which the resistance will double, Let the temperature increase by T °C so \[1.5\times 2=1.5+0.00375\times T\] or \[3=1.5+0.00375\times T\] From this you can fine T, can't you?

OpenStudy (ash2326):

@alfers101 Do you get this?

OpenStudy (anonymous):

yup i am thanks. :D now i can answer this one. thanks

OpenStudy (anonymous):

This is a problem that i want to solve but i can't really answer it. It's a sample problem from my book: The density of a copper is 8.89 is 8.89 g/cm^3 and the ρ=1924μΩ-cm. Determine the length(m) and resistance of 12kg of copper; the cross-section of which in 2.09mm^2.

OpenStudy (ash2326):

@alfers101 Please post this in Physics subject. I'll help you there

OpenStudy (anonymous):

okay thanks :D